OpenStudy (anonymous):

What are the vertices of the hyperbola given by the equation [(y-4)^2/121]-[(x+9)^2/144] = 1?

OpenStudy (anonymous):

Vertices are a distance, 'a' away from the centre. Remember the std eqn: \[(y-k)^2/a^2-(x-h)^2/b^2=1\]

OpenStudy (anonymous):

ok the distance A away so 121 is the vertice?

OpenStudy (anonymous):

no i only have coordinates in the answers.

OpenStudy (anonymous):

121=a^2. You have to take the sqrt. You know the hyperbola goes up & down because its y-x. you also know the center, (h,k). The vertices are 'a' away from the centre. Like, (h,k) +- (0,a)

OpenStudy (anonymous):

k so the center is (9,4) and the vertice is 11 away from the center. but which way?

OpenStudy (anonymous):

here is everything you need

OpenStudy (anonymous):

These are the answers i have . Nothing anyone's told me looks like these :( (-9, 16) and (-9, -8) (-9, 15) and (-9, -7) (2, 4) and (-20, 4) (3, 4) and (-21, 4)

OpenStudy (anonymous):

Do you know if the hyperbola is going up or down?

OpenStudy (anonymous):

no I'm sorry. I may have posted the question wierd. here's just copy and paste What are the vertices of the hyperbola given by the equation open parentheses y minus 4 close parentheses squared over 121 minus open parentheses x plus 9 close parentheses squared over 144 = 1?

OpenStudy (anonymous):

Just give me a moment......

OpenStudy (anonymous):

Sorry, but is this algebra 2, no?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Ohhhh.... ok. At first I thought it was Algebra 1, and right now, I'm completing geometry, so I won't be able to help with Algebra 2 until later this month. I'm so so sorry.

OpenStudy (anonymous):

It's ok. do you know anyone else online that could help though?

OpenStudy (anonymous):

Umm... Hero, Calcmathlete, uri, Loujoelou, ParthKohli.... AccessDenied, and a couple of others....

OpenStudy (anonymous):

u still need help?

OpenStudy (anonymous):

yes please!

OpenStudy (amistre64):

first establish which variable can be zero for the equation to still make sense

OpenStudy (amistre64):

if y=0; -x^2=1 makes no "real" sense; so the hyper opens along parallel to the y axis .... up and down

OpenStudy (amistre64):

the vertex parts can then be evaluated by ignoring the x part and solving for y

OpenStudy (anonymous):

does x still =0 or was that just for the beggining?

OpenStudy (amistre64):

(y-4)^2/121 = 1 (y-4)^2 = 121 y-4 = +- 11 y = 4 +- 11

OpenStudy (amistre64):

the vertexes are only position along the axis assosiated with the y in this case; so the x part is pointless; zero out the x part and we can determine the verts

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

4+11 = 15 4-11 = -7 only one option has those as y part values

OpenStudy (anonymous):

oh so the answer is b!

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

what was RKfitz talking about before?

OpenStudy (amistre64):

prolly something more technical ....

OpenStudy (anonymous):

Ok thank you SO much. <3

OpenStudy (amistre64):

youre welcome, and good luck